OpenStudy (tylermckinney16):

You plan to rent a car for your road trip. If you will be gone for 5 days and plan to drive 650 miles, which of these is the better deal? Show work to explain your answer. Company A: $26 per day; unlimited mileage Company B: $15 per day; $0.14 per mile driven

OpenStudy (tylermckinney16):

@triciaal

OpenStudy (tylermckinney16):

I still dont understand how to do it can u walk me through?

OpenStudy (raffle_snaffle):

What is 15*5 = ?

OpenStudy (raffle_snaffle):

I am walking you through it. Answer my questions.

OpenStudy (tylermckinney16):

75

OpenStudy (tylermckinney16):

@raffle_snaffle

OpenStudy (raffle_snaffle):

Okay, what is 0.14*650 = ?

OpenStudy (tylermckinney16):

91

OpenStudy (tylermckinney16):

@raffle_snaffle

OpenStudy (dmbroussard1978):

91 + 75 = $166 vs 26 × 5 = $130 $26 per day is your better deal it's 36 dollars cheaper

OpenStudy (dmbroussard1978):

Company A

OpenStudy (raffle_snaffle):

@dmbroussard1978 dont give the answer away...